Tom is 4 years older than Jerry.
9 years ago tom was 5x as old as Jerry was then.
How old are they now?

Answer by longjonsilver(2297)

first define 2 variables, based on what you need to find....
Let t = tom's age now
Let j = jerry's age now.
t=j+4 --eqn 1
9 years ago, tom was (t-9) and jerry was (j-9), so t-9 = 5(j-9)
Now solve these 2 equations. Look at my many previous examples of how to do this.
jon